Write the equation in standard form then sketch the graph of the equation
x^2+y^2-4x+2y+5=36

Answer by lwsshak3(11628)

Write the equation in standard form then sketch the graph of the equation
x^2+y^2-4x+2y+5=36
***
complete the square:
x^2-4x+y^2+2y+5=36
(x^2-4x+4)+(y^2+2y+1)+5=36-5+4+1
(x-2)^2+(y+1)^2=36
conic is a circle with radius=6 and center at (2,-1)